OpenStudy (anonymous):

A common inhabitant of human intestines is the bacterium Escherichia coli. A cell of this bacterium in a nutrient-broth medium divides into two cells every 20 minutes. The initial population of a culture is 63 cells. Find the rate of growth after 8 hours. (Round your answer to three decimal places.)

OpenStudy (anonymous):

I know that \[p(t) = 63 e^{2.07944*t}\] And I know that there are 1056964608 cells after 8 hours. This is driving me nuts. I've been trying to figure out what rate of growth means. Relative rate of growth is easy, but this... so frustrated....

OpenStudy (anonymous):

What I mean when I say, "i know that.." from above, is that I figured it out. Although I am certain my math is correct. So the above values are right.

OpenStudy (amistre64):

i got the other one i believe in the end lol

OpenStudy (amistre64):

"medium divides into two cells every 20 minutes. The initial population of a culture is 63" 20min = 1/3 hr 3 ln(2) = r then so after 8 hours that should be: A = 63 e^(8*3ln(2)) = 1.05696461 × 10^9

OpenStudy (amistre64):

thats not the rate of growth tho, thats the population size

OpenStudy (amistre64):

the rate of growth is the derivative of the function when t=8 i believe

OpenStudy (amistre64):

A = 63 e^(3ln(2)*t) A' = 3ln(2)*63 e^(3ln(2)*t)

OpenStudy (amistre64):

there would have been a t' but t'=1 with respect to itself so we tend to toss it out

OpenStudy (amistre64):

2197896113.96054.... http://www.wolframalpha.com/input/?i=3ln%282%29*63+e^%283ln%282%29*8%29

OpenStudy (amistre64):

hopefully that makes sense ....

OpenStudy (anonymous):

No idea what you did here, "20min = 1/3 hr 3 ln(2) = r then so after 8 hours that should be:" I just assumed t was in hours. Made it much simpler.

OpenStudy (anonymous):

If i understand correctly, you did not assume t was in hours, but rather t=20min? If so, how did you get 3 ln(2)=r?

OpenStudy (anonymous):

Nevermind ^^. I realized you weren't calculating rate of growth.

OpenStudy (amistre64):

i tapped into a trick I recall that continuous things have a doubling effect every ln(2)/k where k is either the rate of the time depending on which one you need, ln(2)/r = time ln(2)/t = rate since 1/3 of an hour is the time it takes to double; the rate itself is 3 ln(2)

OpenStudy (amistre64):

the rate OR the time ... accursed typos

OpenStudy (anonymous):

I think you're approaching this from a different way than I am supposed to. It might eventually lead to the same conclusion, but.. here let me paste the ENTIRE question.

OpenStudy (amistre64):

it might be (rt)' in the derivative ... but im pretty sure since its a constant that it kinda takes care of itself. k

OpenStudy (anonymous):

There were 4 parts to this question. I got 3 of them correct. (a) Find the relative growth rate. (Assume t is measured in hours.) (b) Find an expression for the number of cells after t hours. (c) Find the number of cells after 8 hours (d) Find the rate of growth after 8 hours. (Round your answer to three decimal places.) (e) When will the population reach 20,000 cells? (Round your answer to two decimal places.) D is the only one I cannot get. I don't think we are supposed to use what you referred to as 'the doubling rule'. Which i think is just the 'Rule of 72' : ln(2)/ln(1+r)

OpenStudy (anonymous):

Don't work out a, b, c, e above by the way. Web assign said I got them right so no need for checking.. And i meant to say there are 5 parts to this question, i got 4 correct. heh

OpenStudy (amistre64):

2P = Pe^(rt) 2P/P = e^(rt) 2 = e^(rt) ln(2) = ln(e^(rt)) ln(2) = rt ln(2)/t = r t, in hours is 1/3 ln(2)/(1/3) = r = 3 ln(2)

OpenStudy (amistre64):

thats just the algebra

OpenStudy (amistre64):

as you can see, the doubling has nothing to do with the initial size; it is only affected by time

OpenStudy (amistre64):

or how fast, the rate ....

OpenStudy (amistre64):

Find the rate of growth after 8 hours. How fast is the population growing after 8 hours; thats a rate of change question that says we take the derivative

OpenStudy (anonymous):

Okay, see what you did. "t, in hours is 1/3 ln(2)/(1/3) = r = 3 ln(2)" the question D is asking for after 8 hours. So, i plug in 8 rather than 1/3 for t?

OpenStudy (amistre64):

no, not that i am interpreting it. The initial equation for the growth tells us size for any given t in hours. teh rate at which that growth occurs is the dervative of the size

OpenStudy (amistre64):

the size is: \[A(x) = 63e^{3ln(2)\ t}\] \[A'(x) = [3ln(2)\ t]'*63e^{3ln(2)\ t}\] \[A'(x) = 3ln(2)*63e^{3ln(2)\ t}\]

OpenStudy (amistre64):

the rate of growth is then A'(x)

OpenStudy (amistre64):

lol ... i used an x :) A(t) and A'(t)

OpenStudy (amistre64):

if you plug in an 8 instead of the 1/3 your implying that this thing doubles every 8 hours which is false

OpenStudy (anonymous):

Okay yeah, i got that. Look above at the second comment. I just wrote it differently. Instead of writing 3ln(2), i just had 2.0794415. So all I needed to do was find p'(t) from the beginning? God dammit!

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

Thank you kind sir!

OpenStudy (amistre64):

youre welcome, hope it helps ;)

OpenStudy (anonymous):

Sigh, I'm having trouble taking the derivative.. Mind telling me where I'm screwing up in my steps? \[A(x)=63e^{3\ln(2) t}\] *product rule* \[A'(x)= [63 * e^{3\ln(2)t}] + [e^{3\ln(2)t} *0]\] So it comes back to the exact same thing? Since d/dx e^x = e^x???... But I know this can't be right?

OpenStudy (anonymous):

Oh wait, did you do chain rule? How do you use chain rule with e^3ln(2)t? Did you just bring down the entire exponents and take the derivative of them? Weird...

OpenStudy (amistre64):

you learnt the rule for "e" a little off. Every function pops out a chain rule. think of e^x as e^(f(x)) such the derivative of e^(f(x)) = e^(f(x))* f'(x)

OpenStudy (amistre64):

derivatives always pop out an x'; but x' with respect to x = 1 so it tends to be discarded with no afterthought

OpenStudy (anonymous):

d/dx [3ln(2)t = 3ln(2) according to your work above? But shouldn't it just be 3/2. Since d/dx ln(x) = 1/x

OpenStudy (amistre64):

3ln(2) is a constant; there are no variables in it to operate upon. think of it as [Ct]' ; which is just C

OpenStudy (amistre64):

d\dx ln(4) = 0 since ln(4) is just a constant. When something doesnt change, its rate of change is 0

OpenStudy (amistre64):

if your always 5 years old, you never grow up :)

OpenStudy (anonymous):

Silly mistake, i get it. TY

OpenStudy (amistre64):

3ln(2) = 2.079 ...; the derivative of 2.079... is just 0

OpenStudy (amistre64):

;)

OpenStudy (anonymous):

Hmmm, I just plugging in 8 for t into A't = , got 2197896114, and I'm getting that that is wrong. The answer is in format: [answer box here] billion cells per hour. So do i divide by 8.. Don't know what to do.

OpenStudy (amistre64):

2 197 896 114 that is about 2.2 billion